# Tessellated phase diagram and extended criticality in driven quasicrystals and quantum Hall matter

### v1

@@ Abstract
Quasiperiodic systems host a rich array of physical phenomena ranging from localization [1, 2, 3, 4] to fractal spectra [5, 6, 2] to criticality [7, 8, 9] to non-trivial topology [10, 11, 12, 13, 14]. Still richer possibilities arise in the presence of external driving, including a competition between thermalization and many-body localization [15, 16] and potentially extended critical phases featuring multifractal wavefunctions [17, 18, 19, 16, 20, 21, 22, 23, 24, 25, 25, 26, 27, 28, 29]. These phenomena arise from the long-range spatial correlation of quasiperiodicity, which itself is the result of projecting a higher-dimensional periodic structure into the lower-dimensional physical space  [30, 10, 11, 12, 13, 14]. The connection to the higher-dimensional space also gives rise to a degree of freedom called a phason mode. A recent experiment [31] illustrated a connection between two seemingly unrelated localization phenomena by mapping a rapidly oscillating phasonic modulation to linearly polarized monochromatic light in the higher-dimensional space. It is then natural to investigate phenomena induced by arbitrary polarizations of irradiation in the higher-dimensional space.
@@ II.1 Tessellated phase diagram
$=$ $-J_{\mathrm{eff}}\sum_{j}\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+\mathrm{h.c.}+\Delta_{\mathrm{eff}}\cos\left(2\pi\beta j+\kappa\right)\hat{n}_{j},$
@@ II.3 Floquet engineering of multifractality without fine-tuning
In this section, we will show that these control capabilities afforded by the DDAAH model (2) allow preparation of an extended critical phase of matter hosting multifractal wavefunctions without fine-tuning. The possibility of an extended critical phase is of interest from both theoretical and experimental perspectives: usually, multifractal wave functions only appear when a system is finely tuned to a phase boundary or mobility edge, which hinders experimental study.
@@ III From driven Harper-Hofstadter to the generalized Su-Schrieffer-Heeger model
Figure 8: Edge modes of the commensurate DDAAH model under elliptically polarized modulation driven at $\varphi_{0}\approx 2.4048$, the first zero of $\mathcal{J}_{0}(\varphi_{0})$. (a) Quasienergy spectrum for $\beta=1/2$ and $\lambda=1$. Two nearly degenerate edge modes between $0<\kappa<\pi$ are visible around $\varepsilon/J=0$. (b) Quasienergy spectrum for $\beta=1/4$ at $\lambda\approx 6.1047$ (so $|J_{1}/\mathcal{J}_{0}(K_{0})|=1$). A pair of nearly degenerate edge modes are visible between $20<\kappa<\pi/$ 2 and $\pi<\kappa<3\pi/2$ inside the gaps around $\varepsilon/J=0$, and two other quantum-Hall type edge modes also appear. Both calculations are done at $\hbar\omega=15J$, $L=100$ and $K_{0}=1.5$ using the Floquet block Hamiltonian (Eq. 14) with open boundary conditions.
@@ IV Experimental realization
Figure 9: Time evolution of the rms width for the DDAAH model (2) at different locations in the phase diagram. Parameters, the polarization of the modulations, the fitted values of the exponent $\gamma$ and the corresponding phases are shown at the top panel. The time-evolution is shown at the bottom panel, where dotted black lines represent the fit for the exponent $\gamma$. The system is driven for $2000$ cycles at $\hbar\omega=15J$ with $\varphi_{0}=K_{0}$ and data in each case are averaged over 10 randomly selected values of $\kappa$
@@ Appendix A Floquet analysis
$\hat{H}_{m}=\sum_{j=1}^{L}$ $-Je^{im\theta_{x}}\mathcal{J}_{m}(K_{0})\left((-1)^{m}\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+\hat{c}_{j}^{\dagger}\hat{c}_{j+1}\right)$
$+\Delta e^{im\theta_{y}}\mathcal{J}_{m}(\varphi_{0})i^{m}\cos\left(2\pi\beta j+\kappa-\frac{m\pi}{2}\right)\hat{n}_{j}.$
@@ Appendix B High-frequency expansion of modulations with a generic polarization
$\hat{H}^{(1)}=\frac{4J\Delta}{\hbar\omega}\sin(\pi\beta)\sum_{m=1}^{\infty}\frac{i^{m+1}}{m}\sin\left(m(\theta_{y}-\theta_{x})\right)\mathcal{J}_{m}(K_{0})\mathcal{J}_{m}(\varphi_{0})\sum_{j=1}^{L-1}\sin\left(2\pi\beta j+\kappa+\pi\beta-\frac{m\pi}{2}\right)\left(-\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+(-1)^{m}\hat{c}_{j}^{\dagger}\hat{c}_{j+1}\right)$

### v2

@@ Abstract
Quasiperiodic systems host a rich array of physical phenomena ranging from localization [1, 2, 3, 4] to fractal spectra [5, 6, 2] to criticality [7, 8, 9, 10, 11] to non-trivial topology [12, 13, 14, 15, 16]. Still richer possibilities arise in the presence of external driving, including a competition between thermalization and many-body localization [17, 18] and potentially extended critical phases featuring multifractal wavefunctions [19, 20, 21, 18, 22, 23, 24, 25, 26, 27, 27, 11, 28, 29, 30]. These phenomena arise from the long-range spatial correlation of quasiperiodicity, which itself is the result of projecting a higher-dimensional periodic structure into the lower-dimensional physical space  [31, 12, 13, 14, 15, 16]. The connection to the higher-dimensional space also gives rise to a degree of freedom called a phason mode. A recent experiment [32] illustrated a connection between two seemingly unrelated localization phenomena by mapping a rapidly oscillating phasonic modulation to linearly polarized monochromatic light in the higher-dimensional space. It is then natural to investigate phenomena induced by arbitrary polarizations of irradiation in the higher-dimensional space.
@@ II.1 Tessellated phase diagram
$=$ $\sum_{j}-J_{\mathrm{eff}}\left(\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+\mathrm{h.c.}\right)+\Delta_{\mathrm{eff}}\cos\left(2\pi\beta j+\kappa\right)\hat{n}_{j},$
@@ II.3 Floquet engineering of multifractality without fine-tuning
In this section, we will show that these control capabilities afforded by the DDAAH model (2) allow preparation of an extended critical phase of matter hosting multifractal wavefunctions without fine-tuning [8, 10, 11]. The possibility of an extended critical phase is of interest from both theoretical and experimental perspectives: usually, multifractal wave functions only appear when a system is finely tuned to a phase boundary or mobility edge, which hinders experimental study.
@@ III From driven Harper-Hofstadter to the generalized Su-Schrieffer-Heeger model
Figure 8: Edge modes of the commensurate DDAAH model under elliptically polarized modulation ($\delta\theta=-\pi/2$) driven at $\varphi_{0}\approx 2.4048$, the first zero of $\mathcal{J}_{0}(\varphi_{0})$. (a) Quasienergy spectrum for $\beta=1/2$ and $\lambda=1$. Two nearly degenerate edge modes between $0<\kappa<\pi$ are visible around $\varepsilon/J=0$. (b) Quasienergy spectrum for $\beta=1/4$ at $\lambda\approx 6.1047$ (so $|J_{1}/\mathcal{J}_{0}(K_{0})|=1$). A pair of nearly degenerate edge modes are visible between $20<\kappa<\pi/$ 2 and $\pi<\kappa<3\pi/2$ inside the gaps around $\varepsilon/J=0$, and two other quantum-Hall type edge modes also appear. Both calculations are done at $\hbar\omega=15J$, $L=100$ and $K_{0}=1.5$ using the Floquet block Hamiltonian (Eq. 14) with open boundary conditions.
@@ IV Experimental realization
Figure 9: Time evolution of the rms width for the DDAAH model (2) at different locations in the phase diagram. Parameters, the polarization of the modulations, the fitted values of the exponent $\gamma$ and the corresponding phases are shown at the top panel. The time-evolution is shown at the bottom panel, where dotted black lines represent the fit for the exponent $\gamma$. The system is driven for $2000$ cycles at $\hbar\omega=15J$ with $\varphi_{0}=K_{0}$ and data in each case are averaged over 10 randomly selected values of $\kappa$. The initial state is a single-site excitation $\psi(t)=\delta_{j,0}$ and the Hamiltonian (2) is numerically integrated with the second order split-operator method with a time step of $dt\approx 1.26\times 10^{-5}\hbar/J$.
@@ Appendix A Floquet analysis
$\hat{H}_{m}=\sum_{j=1}^{L-1}$ $-Je^{im\theta_{x}}\mathcal{J}_{m}(K_{0})\left((-1)^{m}\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+\hat{c}_{j}^{\dagger}\hat{c}_{j+1}\right)$
$+\sum_{j=1}^{L}\Delta e^{im\theta_{y}}\mathcal{J}_{m}(\varphi_{0})i^{m}\cos\left(2\pi\beta j+\kappa-\frac{m\pi}{2}\right)\hat{n}_{j}.$
@@ Appendix B High-frequency expansion of modulations with a generic polarization
$\hat{H}^{(1)}=\frac{4J\Delta}{\hbar\omega}\sin(\pi\beta)\sum_{m=1}^{\infty}\frac{i^{m+1}}{m}\sin\left(m(\theta_{y}-\theta_{x})\right)\mathcal{J}_{m}(K_{0})\mathcal{J}_{m}(\varphi_{0})\\ \times\sum_{j=1}^{L-1}\sin\left(2\pi\beta j+\kappa+\pi\beta-\frac{m\pi}{2}\right)\left(-\hat{c}_{j+1}^{\dagger}\hat{c}_{j}+(-1)^{m}\hat{c}_{j}^{\dagger}\hat{c}_{j+1}\right).$